# Control Pneumatic Soft Bending Actuator with Feedforward Hysteresis Compensation by Pneumatic Physical Reservoir Computing

### v1

@@ Abstract
The nonlinearities of soft robots bring control challenges like hysteresis but also provide them with computational capacities. This paper introduces a fuzzy pneumatic physical reservoir computing (FPRC) model for feedforward hysteresis compensation in motion tracking control of soft actuators. Our method utilizes a pneumatic bending actuator as a physical reservoir with nonlinear computing capacities to control another pneumatic bending actuator. The FPRC model employs a Takagi-Sugeno (T-S) fuzzy model to process outputs from the physical reservoir. In comparative evaluations, the FPRC model shows equivalent training performance to an Echo State Network (ESN) model, whereas it exhibits better test accuracies with significantly reduced execution time. Experiments validate the proposed FPRC model’s effectiveness in controlling the bending motion of the pneumatic soft actuator with open and closed-loop control systems. The proposed FPRC model’s robustness against environmental disturbances has also been experimentally verified. To the authors’ knowledge, this is the first implementation of a physical system in the feedforward hysteresis compensation model for controlling soft actuators. This study is expected to advance physical reservoir computing in nonlinear control applications and extend the feedforward hysteresis compensation methods for controlling soft actuators.
@@ I INTRODUCTION
Soft robotics is swiftly developing. Pneumatic actuation is a widely adopted driving method in soft robotics. Pneumatic Artificial Muscles (PAMs), with the renowned McKibben design, can produce large force outputs with low self-weights. While PAMs are typically designed to execute linear motions, recent advancements like incorporating physical constraints into the standard PAM structure [1] have enabled PAM-based pneumatic soft actuators to perform bending movements.
@@
Their flexible materials offer soft robots benefits like user-friendliness but also bring them significant control challenges [2], particularly when precision and rapid response are required. One of the problems is hysteresis, which is a time-dependent nonlinearity caused by the Coulomb friction between mechanical components and the viscoelasticity of the soft materials [3]. Hysteresis causes nonlinear loops with dead zones in the actuator’s input-to-output relationship [1], leading to difficulties in precisely controlling their motions.
@@
As a result, one popular method to control soft actuators is integrating a feedforward hysteresis compensation model into a feedback control loop [4, 5]. The feedforward hysteresis compensation models are often built by the combination of a group of mathematical play-operators [4] or Recurrent Neural Networks (RNNs) [5]. RNN-based modelings can capture complex time-dependent dynamics and feature easy implementation. However, neural networks typically have poor interpretability and demand ample data for training. Meanwhile, their gradient-based training approaches can be computationally expensive. Compared with RNNs, hysteresis models based on mathematical play-operators possess better transparency and reasonable physical interpretability [6]. Nonetheless, their development often depends on complex and time-consuming optimization methods, such as genetic algorithms (GA) and particle swarm optimization (PSO) [7].
@@
In light of RNNs’ computational complexity, Reservoir Computing (RC) [8], as a variant of RNNs, has been introduced to predict complex nonlinear behaviors. In RC models like the widely used Echo State Network (ESN), the input signal is projected into a high-dimensional reservoir state via an input layer. The state is updated by a large and randomly initialized network named the ‘reservoir.’ Notably, the input layer and the reservoir are fixed during training, with only the output layer, which handles the reservoir state by linear weights, being trained via linear regression. As it requires no adjustment, the reservoir can be replaced by physical systems like soft bodies [2], thus giving rise to Physical Reservoir Computing (PRC). Since physical systems’ state update does not involve the large ‘reservoir’ used in ESNs—which may contain tens to thousands of nodes [9]—PRC models can show greatly reduced computational demands in their execution, making them suitable for real-time control applications.
@@
PRCs can be used with open-loop and closed-loop forms. Prior research has employed the closed-loop PRC to generate autonomous behaviors of soft robots (or actuators) [10, 11] and the open-loop PRC for sensory feedback in control systems [12], as well as other real-time computational tasks [2]. However, the application of open-loop PRC in feedforward hysteresis compensation models for controlling soft actuators remains unexplored. Furthermore, the PRC, which relies on physical systems with actual existences, may be susceptible to real-world disturbances. While the robustness of closed-loop PRC under such conditions has been examined in [11], the impact of environmental disturbances on the performance of open-loop PRC has not been analyzed in [2] and [12].
@@
The Takagi-Sugeno (T-S) fuzzy model [13] is powerful in nonlinear system identification [14]. It has been employed for precise hysteresis modeling with the nonlinear autoregressive moving average (NARMA) model [14]. The T-S fuzzy model has also been integrated into a multi-ESN framework for enhancing the performance in identifying nonlinear systems [15]. Hence, the T-S fuzzy model is expected to boost the PRC’s capabilities in modeling complex nonlinear dynamics.
@@
This paper proposes a novel approach using PRC in a feedforward hysteresis compensation model to control a soft bending actuator. Our model employs a dual-PAM actuator with two air chambers as a pneumatic physical reservoir. A T-S fuzzy model processes the reservoir’s readout and generates the feedforward hysteresis compensation model’s output. Our proposed method shows better test performance than an ESN model with significantly reduced execution time. Experimental results confirm the proposed model’s effectiveness in open and closed-loop bending motion tracking control. The proposed model’s strong robustness against real-world disturbances has also been experimentally verified.
@@
The following unfolds as Section 2 exhibits the design and hysteresis of the pneumatic actuator. Section 3 describes the feedforward hysteresis compensation models. Section 4 presents the experimental results, and Section 5 concludes.
@@ II-A Structural Design
Fig. 1 (a) shows the pneumatic soft bending actuator to be controlled in this work. Its structure comprises a metallic base, a rubber bladder, a metal end, a linking metal plate, and an inextensible fabric shield. Upon pressurization, the rubber bladder expands radially and contracts axially, following the mechanism of PAMs. The metal plate restricts the actuator’s uniform axial contraction, thus inducing its bending movement, Fig. 1 (b) exhibits. Upon depressurization, the actuator returns to the straight shape, as depicted in Fig. 1 (c).
@@
where $\mathcal{F}_{r}(\cdot)$ is the nonlinear mapping with memory capacity (which is discussed in the next section) that governs the transformation from $P_{i}$ to $P_{o}$ via the dual-PAM actuator’s high-dimensional structure. The passive PAM’s inner pressure $P_{o}$ is used as the readout of the physical reservoir’s state.
@@ II-B Hysteresis Characteristics
We used the experimental setup shown in Fig. 3 to obtain the hysteresis characteristics of the soft bending actuator and the physical reservoir. This setup comprises a pneumatic bending actuator and a pneumatic physical reservoir (both provided by Bridgestone), two proportional valves (Festo MPYE-5-M5-010B), an angular sensor (BENDLABS-1AXIS), three pressure sensors (SMC PSE540A-R06), and a computer with a Contec AI-1616L-LPE analog input board and a Contec AO-1608L-LPE analog output board. All experiment signal refresh rates were 200 Hz in this work.
@@
Two hysteresis experiments were conducted by pressurizing the bending actuator with different inputs $P_{\text{exp}}$ [kPa]: the first applied a sweep-frequency input shown in Fig. 4 (a), and the second used a complex input shown in Fig. 4 (c). During the experiments, the actuator’s bending angle $\theta$ [deg] was converted into $P_{i}$ by $P_{i}(k)=K_{\text{in}}\theta(k)$. $K_{\text{in}}=7.0$ was fixed throughout this study; this value was determined based on the physical reservoir’s actuation upper limit of 450 [kPa] and the bending actuator’s output range of approximately 60 [deg]. We used the bending angle $\theta$ to generate $P_{i}$ at this stage to obtain the response of $P_{o}$ corresponding to $\theta$, facilitating the subsequent inverse operation in the development of the feedforward model. The pressure $P_{i}$ was regulated by a proportional valve and a feedback proportional-integral (PI) controller, whose gain settings are detailed in Section IV.
@@
Figure 3: Experimental apparatus. Blue and black lines indicate the pneumatic and electronic circuits, respectively.
@@
The hysteresis between $\theta$ and $P_{\text{exp}}$ obtained from the first and second experiments are depicted in Fig. 4 (b) and (d), respectively. Similarly, the hysteresis between $P_{o}$ and $\theta$ during these experiments are shown in Fig. 5 (a) and (b). The hysteresis loops in both Fig. 4 and Fig. 5 are frequency-independent, since the primary cause of the PAM-based soft actuators’ hysteresis is the rate-independent Coulomb friction [3]. As a time-dependent nonlinearity, the hysteresis observed in Fig. 4 hinders the soft actuator’s precise motion control [4], while that shown in Fig. 5 realize the physical reservoir’s nonlinear response to input signals and its memory capacities (caused by the time-dependent properties of hysteresis).
@@
Figure 4: Hysteresis experiments: (a) and (b) Sweep-frequency pressure input and corresponding hysteresis loops of the first experiment; (c) and (d) Complex pressure input and hysteresis loops of the second experiment. The signal in (a) lasts 120 [s] and sweeps the frequency range from 0.1 Hz to 1 Hz. The signal in (c) is $17535\sum_{i=1}^{5}\sin\left(2\pi f_{i}t-\pi/2\right)+$ 175 [kPa], $t\in[0,80]$ is the time [s], $f_{1}=0.12$, $f_{2}=0.04$, $f_{3}=0.31$, $f_{4}=0.29$, and $f_{5}=0.25$.
@@
Figure 5: Hysteresis experiments: (a) Hysteresis between $P_{o}$ and $\theta$ of the pneumatic actuator during the first experiment; (b) The corresponding hysteresis relationship obtained from the second experiment.
@@ III Feedforward Hysteresis Compensation
In this section, we build two different feedforward hysteresis compensation models and conduct comparisons. The first model is based on a conventional ESN, and the second is the T-S fuzzy PRC (FPRC) model proposed in this work.
III-A ESN Hysteresis Compensation Model
Fig. 6 is the schematic of the control system using the ESN feedforward model to control the soft actuator’s bending motion. The system has a closed feedback loop and an open feedforward loop, where $e_{\theta}=\theta_{d}-\theta$ [deg] is the bending motion tracking control error, $P_{\text{fb}}$ [kPa] is the feedback adjustment from the proportional-differential (PD) controller in the closed loop, and $P_{\text{ff}}$ [kPa] is the feedforward hysteresis compensation model’s output. The desired actuation pressure $P_{d}=P_{\text{ff}}+P_{\text{fb}}$ [kPa], is regulated by a PI pressure controller and applied to the soft bending actuator. The PI pressure controller’s performance is assumed to be ideal in this study.
@@ III-A ESN Hysteresis Compensation Model
$\mathbf{x}(k+1)=\gamma\ \mathbf{x}(k)+(1-\gamma)\tanh\left(\mathbf{w}_{\text{input}}\theta_{d}(k)+\mathbf{W}_{r}\mathbf{x}(k)\right),$ (3)
where $\gamma\in[0,1)$ is the leaky rate, and $\tanh(\cdot)$ is the hyperbolic tangent activation function commonly used in ESN models. The hidden state $\mathbf{x}$ is combined with the reference signal’s temporal sequence $\bm{\theta}(k)=[\theta_{d}(k),\theta_{d}(k-1),\cdots,\theta_{d}(k-n_{y}+1)]^{\top}\in\mathbb{R}^{n_{y}}$ and a unit element to form the extended state $\mathbf{x}^{*}(k)=[1,\bm{\theta}^{\top}(k),\mathbf{x}^{\top}(k)]^{\top}\in\mathbb{R}^{r+n_{y}+1}$. The extended state $\mathbf{x}^{*}$ is processed by the output layer $\mathbf{w}_{\text{out}}\in\mathbb{R}^{r+n_{y}+1}$ to produce:
@@
where $P_{\text{ff}}$ is the model’s output corresponding to $\theta_{d}$. $\mathbf{w}_{\text{input}}$ and $\mathbf{W}_{r}$ are initially randomized with a $(0,1)$ distribution and subsequently fixed after being adjusted by the scaling factor $f_{\text{input}}$, which scales the input weights in $\mathbf{w}_{\text{input}}$, and the spectral radius $\rho$, which adjusts the largest absolute eigenvalue of $\mathbf{W}_{r}$. The only trainable parameters in the model are the weights in $\mathbf{w}_{\text{out}}$, which are optimized using linear regression with Tikhonov regularization (ridge regression) as:
@@
III-B Fuzzy PRC Hysteresis Compensation Model
In our proposed FPRC feedforward hysteresis compensation model, the reservoir $\mathbf{W}_{r}$ in ESN is substituted by the dual-PAM actuator shown in Fig. 2. Fig. 7 is a schematic of the control system incorporating the FPRC feedforward model, where the closed-loop components are consistent with those depicted in Fig. 6. The FPRC model uses $K_{\text{in}}$ to convert $\theta_{d}$ into $P_{i}$ following (1). $P_{i}$ is regulated by a PI pressure controller, whose performance is assumed to be ideal.

### v2

@@ Abstract
The nonlinearities of soft robots bring control challenges like hysteresis but also provide them with computational capacities. This paper introduces a fuzzy pneumatic physical reservoir computing (FPRC) model for feedforward hysteresis compensation in motion tracking control of soft actuators. Our method utilizes a pneumatic bending actuator as a physical reservoir with nonlinear computing capacities to control another pneumatic bending actuator. The FPRC model employs a Takagi-Sugeno (T-S) fuzzy logic to process outputs from the physical reservoir. The proposed FPRC model shows equivalent training performance to an Echo State Network (ESN) model, whereas it exhibits better test accuracies with significantly reduced execution time. Experiments validate the FPRC model’s effectiveness in controlling the bending motion of a pneumatic soft actuator with open-loop and closed-loop control system setups. The proposed FPRC model’s robustness against environmental disturbances has also been experimentally verified. To the authors’ knowledge, this is the first implementation of a physical system in the feedforward hysteresis compensation model for controlling soft actuators. This study is expected to advance physical reservoir computing in nonlinear control applications and extend the feedforward hysteresis compensation methods for controlling soft actuators.
Index Terms:
Soft Robotics, Physical Reservoir Computing
@@ I Introduction
Pneumatic actuation is a widely favored driving method in soft robotics. Pneumatic Artificial Muscles (PAMs), with the renowned McKibben design [1], can produce large force with low self-weights. While PAMs are designed to execute linear motions, modifications like incorporating physical constraints into the standard PAM structure [2] have enabled PAM-based soft actuators to perform bending movements.
@@
Their flexible materials offer soft robots benefits like user-friendliness but also bring them significant control challenges [3], particularly when precision and rapid response are required. One of the problems is hysteresis, which is a time-dependent nonlinearity caused by the Coulomb friction between mechanical components and the viscoelasticity of the soft materials [1]. Hysteresis causes nonlinear loops with dead zones in soft actuators’ input-to-output relationships [2], leading to difficulties in precisely controlling their motions.
@@
As a result, one popular method to control soft actuators is integrating a feedforward hysteresis compensation model into a feedback control loop [4, 5]. The feedforward hysteresis compensation models are often built by the combination of a group of mathematical play-operators [4] or using Recurrent Neural Networks (RNNs) [5]. RNN-based modelings can capture complex time-dependent dynamics and feature easy implementation. However, neural networks typically have poor interpretability and require a large amount of data for training. Meanwhile, their gradient-based training approaches can be computationally expensive. Compared with RNNs, hysteresis models based on mathematical play-operators possess better transparency and reasonable physical interpretability [6]. Nonetheless, their development usually depends on complex and time-consuming optimization techniques, such as genetic algorithms (GA) and particle swarm optimization (PSO) [7].
@@
In light of classic RNNs’ computational complexity, Reservoir Computing (RC) [8], as a variant of RNNs, has been introduced to predict complex nonlinear behaviors. In RC models like the widely used Echo State Network (ESN) [8], the input signal is projected into a high-dimensional reservoir state via an input layer. The state is updated by a large and randomly initialized network named the ‘reservoir.’ Notably, the input layer and the reservoir are fixed during training, with only the output layer, which handles the reservoir state by linear weights, being trained with linear regressions. As it requires no adjustment, the reservoir can be replaced by physical systems like soft bodies [3], thus giving rise to physical reservoir computing (PRC). Since physical systems’ state update does not involve the large ‘reservoir’ used in ESNs—which may contain tens to thousands of nodes [9]—PRC models can show greatly reduced computational requirements in their execution [10], making them suitable for real-time control applications.
@@
PRCs can be used with open-loop and closed-loop forms. Prior research has employed the closed-loop PRC to generate autonomous behaviors of soft robots (or actuators) [10, 11] and the open-loop PRC for sensory feedback in control systems [12], as well as other real-time computational tasks [3]. However, the application of open-loop PRC in feedforward hysteresis compensation models for controlling soft actuators remains unexplored. Furthermore, the PRC, which relies on physical systems with actual existences, may be susceptible to real-world disturbances. While the robustness of the closed-loop PRC under such conditions has been examined in [11], the impact of environmental disturbances on the performance of open-loop PRC has not been investigated in [3] and [12].
@@
The Takagi-Sugeno (T-S) fuzzy logic [13] is a powerful tool in nonlinear system identifications [14]. It has been utilized for precise hysteresis modeling with the nonlinear autoregressive moving average (NARMA) model [14]. The T-S fuzzy logic has also been integrated into a multi-ESN model for enhancing the performance in identifying nonlinear systems [15]. Hence, we consider the T-S fuzzy logic also capable of improving the PRC’s capabilities in modeling complex nonlinear dynamics.
@@
This paper proposes a novel control approach using PRC in a feedforward hysteresis compensation model to control a soft bending actuator. Our method employs a dual-PAM actuator with two air chambers as a pneumatic physical reservoir. A T-S fuzzy model is applied to process the reservoir’s readouts and generate the feedforward hysteresis compensation model’s output. The proposed method shows better accuracies than an ESN and significantly reduced execution time in the model test phase. Experimental results confirm the proposed model’s effectiveness in open and closed-loop bending motion tracking control. The proposed model’s strong robustness against real-world disturbances has also been experimentally demonstrated.
@@
The remainder unfolds as follows: Section 2 presents the design and hysteresis of the pneumatic soft actuator; Section 3 describes the feedforward hysteresis compensation models; Section 4 exhibits experimental results; Section 5 concludes.
@@ II-A Structural Design
Fig. 1 shows the pneumatic soft bending actuator, whose motion is this work’s control object. Its structure comprises a metallic base, a rubber bladder, a metal end, a linking metal plate, and an inextensible fabric shield. Upon pressurization, the rubber bladder expands radially and contracts axially, following the mechanism of PAMs. The metal plate restricts the actuator’s uniform axial contraction, thus inducing its bending, as Fig. 1 (b) illustrates. Upon depressurization, the actuator returns to its straight shape, as depicted in Fig. 1 (c).
@@
where $\mathcal{F}_{r}(\cdot)$ denotes the nonlinear mapping with memory capacity (which is discussed in the next section) that governs the transformation from $P_{i}$ to $P_{o}$ through the dual-PAM actuator’s high-dimensional structure. The passive PAM’s inner pressure $P_{o}$ is employed as the readout of the physical reservoir’s state.
@@ II-B Hysteresis Characteristics
We used the experimental setup shown in Fig. 3 to obtain the hysteresis properties of the soft bending actuator and the physical reservoir. This setup comprises the soft actuator and the physical reservoir (both provided by Bridgestone), two proportional valves (Festo MPYE-5-M5-010B), an angular sensor (BENDLABS-1AXIS), three pressure sensors (SMC PSE540A-R06), and a computer with a Contec AI-1616L-LPE analog input board and a Contec AO-1608L-LPE analog output board. All experiment signal refresh rates were 200 Hz.
@@
Two hysteresis experiments were conducted by pressurizing the bending actuator with different experiment pressure inputs $P_{\text{exp}}$ [kPa]: the first used a sweep-frequency signal shown by Fig. 4 (a), and the second used a complex input shown by Fig. 4 (c). During the experiments, the actuator’s bending angle $\theta$ [deg] was fed into the physical reservoir by $P_{i}(k)=K_{\text{in}}\theta(k)$. $K_{\text{in}}=7.0$ was fixed throughout this study, this value was chosen based on the physical reservoir’s actuation limit of 450 [kPa] and the bending actuator’s output range of approximately 60 [deg]. We used the bending angle $\theta$ to generate $P_{i}$ at this stage to obtain the response of $P_{o}$ corresponding to $\theta$, facilitating the subsequent inverse operation in the development of the feedforward model. The pressure $P_{i}$ was regulated by a proportional valve and a feedback proportional-integral (PI) controller, whose gain settings will be detailed in Section IV.
@@
Figure 3: Experimental apparatus. Blue and black lines in the right-located figure indicate the pneumatic and electronic circuits, respectively.
@@
The hysteresis between $\theta$ and $P_{\text{exp}}$ obtained from the first and second experiments are depicted in Fig. 4 (b) and (d), respectively. Similarly, the hysteresis between $P_{o}$ and $\theta$ during the two experiments are presented in Fig. 5 (a) and (b). The hysteresis loops in both Fig. 4 and Fig. 5 are frequency-independent since the primary cause of the PAM-based soft actuators’ hysteresis is the rate-independent Coulomb friction [1]. As a time-dependent nonlinearity, the hysteresis observed in Fig. 4 hinders the soft actuator’s precise motion control [4], while that shown in Fig. 5 realizes the physical reservoir’s nonlinear responses to input signals and its memory capacity, which is caused by the time-dependent properties of hysteresis.
@@
Figure 4: Hysteresis experiments: (a) and (b) Sweep-frequency pressure input $P_{\text{exp}}$ and the corresponding hysteresis loops of the first experiment; (c) and (d) Complex pressure input $P_{\text{exp}}$ and hysteresis loops of the second experiment. The signal in (a) lasts 120 [s] and sweeps the frequencies from 0.1 Hz to 1 Hz. The signal in (c) is $17535\sum_{i=1}^{5}\sin\left(2\pi f_{i}t-\pi/2\right)+$ 175 [kPa], where $t\in[0,80]$ is the time [s], $f_{1}=0.12$, $f_{2}=0.04$, $f_{3}=0.31$, $f_{4}=0.29$, and $f_{5}=0.25$.
@@
Figure 5: Hysteresis experiments: (a) Hysteresis between $P_{o}$ and the soft actuator’s bending $\theta$ obtained from the first hysteresis experiment; (b) The corresponding hysteresis relationship obtained from the second experiment.
@@ III Feedforward Hysteresis Compensation
In this section, we build two different feedforward hysteresis compensation models and conduct comparisons. The first model is based on a conventional ESN, and the second one is the T-S fuzzy PRC (FPRC) model proposed in this work.
III-A ESN Feedforward Hysteresis Compensation Model
Fig. 6 is the schematic of the control system using the ESN feedforward model to control the soft actuator’s bending motion. The system has a closed feedback loop and an open feedforward loop, where $e_{\theta}=\theta_{d}-\theta$ [deg] is the bending motion tracking control error, $P_{\text{fb}}$ [kPa] is the feedback adjustment from the proportional-differential (PD) controller in the closed loop, and $P_{\text{ff}}$ [kPa] is the feedforward hysteresis compensation model’s output. The desired actuation pressure $P_{d}=P_{\text{ff}}+P_{\text{fb}}$ [kPa], is regulated by a PI pressure controller and used to pressurize the soft bending actuator. The PI pressure controller’s performance is assumed to be ideal in this study.
@@ III-A ESN Feedforward Hysteresis Compensation Model
$\mathbf{x}(k+1)=\gamma\ \mathbf{x}(k)+(1-\gamma)\tanh\left(\mathbf{w}_{\text{in}}\theta_{d}(k)+\mathbf{W}_{r}\mathbf{x}(k)\right),$ (3)
where $\gamma\in[0,1)$ is the leaky rate, and $\tanh(\cdot)$ is the hyperbolic tangent activation function commonly used in ESN models. The hidden state $\mathbf{x}$ is combined with the reference signal’s temporal sequence $\bm{\theta}(k)=[\theta_{d}(k),\theta_{d}(k-1),\cdots,\theta_{d}(k-n_{y}+1)]^{\top}\in\mathbb{R}^{n_{y}}$ and a unit element 1 to form the extended state $\mathbf{x}^{*}(k)=[1,\bm{\theta}^{\top}(k),\mathbf{x}^{\top}(k)]^{\top}\in\mathbb{R}^{r+n_{y}+1}$. The extended state $\mathbf{x}^{*}$ is processed by the output layer $\mathbf{w}_{\text{out}}\in\mathbb{R}^{r+n_{y}+1}$ to produce:
@@
where $P_{\text{ff}}$ is the feedforward model’s output corresponding to the given desired bending angle $\theta_{d}$. $\mathbf{w}_{\text{in}}$ and $\mathbf{W}_{r}$ are initially randomized with parameters from a $N(0,1)$ distribution and subsequently fixed after being adjusted by the scaling factor $f_{\text{input}}$, which scales the input weights in $\mathbf{w}_{\text{in}}$, and the spectral radius $\rho$, which adjusts the largest absolute eigenvalue of $\mathbf{W}_{r}$. The only trainable parameters in the model are linear weights in $\mathbf{w}_{\text{out}}$, which are optimized using the linear regression with Tikhonov regularization (ridge regression) as expressed by:
@@
The dataset $K_{\text{train}}$ comprises $P_{\text{exp}}$, $P_{o}$, and $\theta$. The objective of training the feedforward model is to obtain an approximation of the inverse projection depicted in Fig. 4 that can estimate the actuator’s input $P_{\text{exp}}$ given the output $\theta$ (and its nonlinear transformation performed by the ESN or the PRC). Hence, the experimentally obtained $\theta$ serves as the desired $\theta_{d}$ in (1), (3), and (4) in the model training and test phases. The training process seeks to satisfy $P_{\text{ff}}\approx P_{\text{exp}}$ by the optimization shown in (5) with $P_{\text{exp}}$ shown in Fig. 4 (a) and (c) as the ground truth.
III-B Fuzzy PRC Feedforward Hysteresis Compensation Model
Fig. 7 presents a schematic of the control system with the FPRC feedforward model, where the closed-loop components are consistent with Fig. 6 with the reservoir part $\mathbf{W}_{r}$ in the ESN being replaced by the physical reservoir in Fig. 2. The FPRC model uses $K_{\text{in}}$ to convert $\theta_{d}$ into $P_{i}$ by (1). $P_{i}$ is regulated by a PI pressure controller, whose performance is assumed ideal.
It is essential to clarify that while the performance of the PI pressure controllers is deemed ideal in both the closed-loop control system and the FPRC model, these assumptions are based on different reasons. The PI controller is considered ideal in the closed-loop control system as the pressure control task is relatively simple and typically performs satisfactorily compared to controlling the soft actuator’s bending motion. Therefore, the pressure control is not the primary focus of this research. Conversely, the assumption of ideal performance for the PI controller in the FPRC model is based on the fact that the proportional valve, shown in Fig. 3 and controlled by the PI controller, also has its dynamics. The valve’s dynamics and any pressure control errors are integrated into the overall dynamics of the nonlinear physical system in the FPRC model. Thus, the pressure control errors do not require consideration.